A regular six-sided die is rolled 9 times. What is the probability of getting a 1 or 6 on exactly 7 of those rolls?
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Prob ( getting 1 or 6 ) on a 6 sided dice = 2/6 = 1/3

Now when the die is rolled 9 times, we choose 7 rolls out of these

=> 9C7

Now in the remaining rolls we have to get other numbers than 1 ir 6 => 4/6 = 2/3

Hence the total probability

=> 9C7 * (1/3)7* (2/3) * (2/3)

= 0.01646 * 4/9

= 0.0077316 or 144/19683